![]() |
|||||||||||||||
Welcome to Our Site! |
|||||||||||||||
Progress Report > Package III - Electrical Works
Summary from Contractor The month of April saw major progress, as expected, in the testing and pre-commissioning of power, CCTV, telecommunications, fire alarm and detection, and traffic control system. Testing was ongoing while final installations of parts of these systems were being completed. This included OCC control room equipment and workstation connections and use and installation and commissioning of the VDU. Over all commissioning by ABB was at approximately 50% at end of April. All main tunnel and evacuation tunnel lighting was operational from local control by late April. The Contractor continued with installation works that should have been finished months ago at STP and NTP and in VMB. All substations except S/S#8 at STP were energised by end of April. The MV, fibre optic and road lighting cables were installed along the south road after months of delay in slow excavation of the cable trench. On site training began in April with HAMADECO staff assigned to accompany PKIII trainers during field final installations and checking. Installation of road lighting poles and luminaires along the south road and connections was almost completed as was the last road lights from NTP to Highway 1. Final assembly of the Radio Rebroadcast system was ongoing in Belgium and FAT is expected in May. Completion of installation works continues for parts of CCTV, traffic control, telephone, and for general electrical works. PKIII have scheduled 1 month for completion of RRB in June. Main panel hardware and antenna mast has not yet been delivered. Physical Progress in this month Completion of installation and testing of MV cable and fibre optic cable along south road (OCC to STP). The last 400m of MV cable for second feeder to S/S#6 from S/S#3 was installed, tested and energised. This was missing because of Contractors errors in MV cable quantity order. Installation of GIS, distribution transformers, main LV switchboard, UPS, battery systems, and all other major items installation completed in all substations. Traffic control system loop detector installation with installation and connection of feeder cables was for loops inside walkway was completed. Installation of the fire detection fibre optic cable sensor in the MT was completed. Installation of manual push button at fire hydrant niches began and was 90% complete Installation of the radio rebroadcast leaky coaxial antenna in the tunnel continued but is not complete. Installation of CCTV cameras and associated connection boxes was mainly carried out in April but 9 units remained at month’s end. Excavation of trench for MV, LV, and fibre optic cable along the south road continues and was nearing completion at the end of February. Road lighting poles and luminaires 90% completed along north and south roads and bridges. Various cable ladder and small power and lighting installations in tunnel and VAB were being completed. Fire resistant (FR) cable joint kits were delivered and installed in FR cables in the tunnel. This homepage is presented by the Consultant Team Any further question should be forwarded to Webmaster or Consultant Team
|
|
||||||||||||||